On a school trip the ratio of the number of teachers to students is 1:15. The ratio of male students to female students is 7:5. Work out what percentage of people on the trip are female. Give your answer correct to the nearest whole number.
I've tried multiplying by 5 but I just don't get ratios.
Please can anyone help me with this question?

Answer by ankor@dixie-net.com(22740)

On a school trip the ratio of the number of teachers to students is 1:15.
The ratio of male students to female students is 7:5.
Work out what percentage of people on the trip are female.
Give your answer correct to the nearest whole number.
I've tried multiplying by 5 but I just don't get ratios.
:
you have to know the male to female teachers ratio also
Let's assume the male to female teachers was 3:5, that means we have a minimum of 8 teachers
The teacher student ratio is given as 1:15, therefore we have 15 * 8 = 120 students
The total people, students and teachers = 128
We know we have  5 female teachers. Find how many female students
let x = the multiplier, then
7x = number of male students
5x = number of female students
7x + 5x = 120
12x = 120
x = 120/12
x = 10
therefore the number of female students 10*5 = 50 female students
Total female teachers and female students 5 + 50 = 55
Find the percentage females (55)  to total in the group (128)
55/128  * 100 = 43% females in the group